# Temperature Optimization for Bayesian Deep Learning

### v1

@@ Abstract
The Cold Posterior Effect (CPE) is a phenomenon in Bayesian Deep Learning (BDL), where tempering the posterior to a cold temperature often improves the predictive performance of the posterior predictive distribution (PPD). Although the term ‘CPE’ suggests colder temperatures are inherently better, the BDL community increasingly recognizes that this is not always the case. Despite this, there remains no systematic method for finding the optimal temperature beyond grid search. In this work, we propose a data-driven approach to select the temperature that maximizes test log-predictive density, treating the temperature as a model parameter and estimating it directly from the data. We empirically demonstrate that our method performs comparably to grid search, at a fraction of the cost, across both regression and classification tasks. Finally, we highlight the differing perspectives on CPE between the BDL and Generalized Bayes communities: while the former primarily focuses on predictive performance of the PPD, the latter emphasizes calibrated uncertainty and robustness to model misspecification; these distinct objectives lead to different temperature preferences.
@@ 1 Introduction
The last two decades have seen substantial advances in the development of machine learning tools. This has led to deployment across a range of application domains, from image recognition and natural language processing to autonomous vehicles and medical diagnostics, to name a few. Driven in no small part by the technological advances that have enabled large-scale training of high-fidelity models, these models obtain impressive generalization performance and accuracy that are not readily explained with classical statistical wisdom.
@@
There is growing recognition in the BDL community that, despite the term ‘CPE’, colder temperatures do not always result in better predictive performance for the PPD (Adlam et al., 2020; Zhang et al., 2024). Unfortunately, the common approach of temperature tuning via grid search is computationally expensive, as it requires extra posterior sampling for multiple $T$ across the grid [^1] Consider ResNet20-CIFAR10 for example. Our approach requires only 35 minutes to find the optimal $T$, and an additional 327 minutes to compute the posterior at that optimal $T$. In contrast, a grid search across 10 temperatures takes a substantial 3270 minutes, resulting in a 9x speedup with our method.. Sampling from a cold posterior is challenging and requires careful tuning of the sampler with temperature diagnostics (Wenzel et al., 2020). To address these issues, we propose a data-driven method to select an appropriate temperature. To the best of our knowledge, no such dedicated tool exists without appealing to intermediate approximations (e.g., variational inference (Laves et al., 2021)). Our method only requires maximizing a likelihood function to find a suitable temperature. This can usually be done as part of the sampler warm-up phase, and importantly, does not require any extra posterior sampling.
@@ 2 Tempered Posteriors and the Cold Posterior Effect
$p_{\beta}(y|x,\mathcal{D})\coloneq\mathbb{E}_{p_{\beta}(\theta|\mathcal{D})}[p(y|x,\theta)]=\int p(y|x,\theta)p_{\beta}(\theta|\mathcal{D})\,\textrm{d}\theta,$ (3)
which is constructed from Bayesian model averaging, can achieve better performance (in terms of the test LPD defined below) in regression and classification by artificially tempering the posterior to $\beta>1$ (Wenzel et al., 2020). While our construction in (2) follows the convention of tempering both the likelihood and the prior (Wenzel et al., 2020; Fortuin et al., 2022), improvements in performance have also been observed when only tempering the likelihood (Aitchison, 2021; Bachmann et al., 2022; Kapoor et al., 2022).
@@
where the hat over $p$ indicates dependence on the training data $\mathcal{D}$, and the expectation is only taken over the ‘new’ observation $(x,y)$. Note that the test LPD is sometimes referred to as the ‘test log-likelihood’ in the literature, and the negative test LPD is often called ‘negative log-likelihood (NLL)’. However, we will avoid these terms as they can sound ambiguous. It can be shown that a predictive density $\hat{p}(y|x)$ with a higher test LPD is closer to the truth $q(y|x)$ in the sense that the Kullback-Leibler (KL) divergence $D_{\mathrm{KL}}(q(y|x)\|\hat{p}(y|x))$ is smaller. Our primary interest in this work is to select $\beta$ such that the test LPD of (3) is high.
@@ 2.1 Outline
Following the reasoning that the aleatoric uncertainty in the data can be quantified by $\beta$ (Adlam et al., 2020; Kapoor et al., 2022), we advocate a likelihood-based approach to select $\beta$ directly from the data; details are presented in Section 3. We empirically show that our method can select a $\beta$ that is near-optimal according to the LPD metric. Experimental results are presented in Section 4.
We then conclude with a detailed discussion aimed at dispelling some common misconceptions surrounding the CPE. In particular, we review recent work in the BDL community that explains why colder temperatures do not always lead to better performance in terms of the test LPD metric. We also address the converse hypothesis sometimes encountered outside the BDL literature—that warmer temperatures are often better. This can hold, but only for metrics that differ fundamentally from test LPD. Specifically, in Sections 5 and 6, we examine existing literature on CPE from the perspectives of the BDL and Generalized Bayes (GB) communities, respectively. We emphasize that these two communities prioritize different objectives: in BDL, the primary focus is on maximizing the predictive performance of the PPD, while in GB, the emphasis lies on robustness to model misspecification and calibrated uncertainty. Furthermore, the types of statistical models and datasets explored in each community are notably distinct. In BDL, large-scale neural networks are trained on extensive, nearly noiseless datasets, whereas in GB, the datasets are typically smaller, noisier, and the models are classical regular statistical models. Consequently, it is not surprising that each community arrives at different recommendations regarding temperature tuning.
@@ 3 Temperature Selection for Test LPD
It is commonly believed that the PPD (3) at $\beta=1$ is optimal for test LPD when the model is well-specified (Adlam et al., 2020; Aitchison, 2021). However, this is not always the case (see Appendix A for a counter-example). In fact, Zhang et al. (2024) argues that in the likelihood tempering case, $\beta=1$ can only be optimal if the expected Gibbs training loss remains unchanged with the inclusion of new data. We follow the reasoning in Zhang et al. (2024) and reach a similar conclusion for posterior tempering; see Appendix B for details. This motivates the development of an efficient method to select the optimal $\beta$.
If we could have some theoretical grasp on how temperature affects the test LPD of the PPD in (3), that is, $\operatorname{LPD}(p_{\beta}(y|x,\mathcal{D}))$, then it might suggest a methodology for temperature selection. A little known result to the BDL and GB communities is the following from singular learning theory (Watanabe, 2010a, Lemma 3), which applies to both regular and singular models $p(y|x,\theta)$:
$-\mathbb{E}_{\mathcal{D}}\operatorname{LPD}(p_{\beta}(y|x,\mathcal{D}))$ $=-\mathbb{E}_{\mathcal{D}}\mathbb{E}_{q(x,y)}\log p_{\beta}(y|x,\mathcal{D})$
$=-\mathbb{E}_{q(x,y)}\log p(y|x,\theta_{\dagger})+\left[\frac{\lambda-\nu(\beta)}{\beta}+\nu(\beta)\right]\frac{1}{n}+o\left(\frac{1}{n}\right)$ (4)
@@
Some comments on (4) are in order. This result allows for both misspecification ($\theta_{\dagger}$ is not necessarily such that the KL divergence between the truth and $p(y|x,\theta_{\dagger})$ is zero) and singular models singular models $p(y|x,\theta)$, including neural networks, as well as classical models satisfying standard regularity conditions. On the other hand, it is an asymptotic result in the sample size $n$, and hence the prior plays no role. We also remark that the relation in (4) pertains to the average (negative) test LPD of the PPD in (3), with the average taken over the training set $\mathcal{D}$, as indicated by the notation $\mathbb{E}_{\mathcal{D}}$. More precise interpretation of (4) can be divided into two settings:
@@
For well-specified regular models, $\lambda=\nu(\beta)=d/2$ for all $\beta$ (Watanabe, 2009), and the second term, $\left[\frac{\lambda-\nu(\beta)}{\beta}+\nu(\beta)\right]\frac{1}{n}$, reduces to $\frac{d}{2n}$. This implies that for large $n$, temperature has little impact (though it may appear in higher-order terms in the expansion). This aligns with recent work by McLatchie et al. (2024), which arrives at a similar conclusion using different techniques. Under model misspecification, the second term can alternatively be expressed as $\frac{\beta}{n}\mathbb{E}_{\mathcal{D}}V(n)$ (Watanabe, 2010b), where $V(n)$ represents the functional variance and now depends on the temperature in a non-trivial manner.
@@ 3.1 Selecting Temperature using the Tempered Model
In this section, we introduce a method to select $\beta$ that corresponds to high test LPD in both regression and classification tasks. Our approach is grounded in the insight that the tempered posterior can be reframed as the posterior from an alternative model-prior pair (Zeno et al., 2020; Zhang et al., 2024). Following these works, we define the tempered model,
@@
where the ‘rest of the terms’, $\tilde{p}(\theta|\beta)\propto p(\theta)^{\beta}\prod_{x\in\mathcal{D}}\int p(y^{\prime}|x,\theta)^{\beta}\textrm{d}y^{\prime}$, can be viewed as a prior on $\theta$ with the ‘normalizing constant’ being a function of $x$. Note that $\tilde{p}(\theta|\beta)$ can be seen as an input-dependent prior (Zeno et al., 2020) but we suppress the dependence on $x$ in the notation. The (inverse) temperature $\beta$ has an intuitive interpretation of controlling the ‘spikiness’ of the tempered model. We give two examples with Gaussian (regression) and categorical (classification) models:
@@
$p(y|x,\theta,\beta)=\mathcal{N}\left(y\left|\,\mu(x;\theta),\frac{\sigma^{2}}{\beta}\right.\right),\quad\tilde{p}(\theta|\beta)\propto p(\theta)^{\beta}\left(\dfrac{2\pi\sigma^{2}}{\beta(2\pi\sigma^{2})^{\beta}}\right)^{n/2}.$
@@
$p(y|x,\theta,\beta)=\dfrac{\exp(\beta f_{y}(x;\theta))}{\sum^{K}_{k=1}\exp(\beta f_{k}(x;\theta))},\quad\tilde{p}(\theta|\beta)\propto p(\theta)^{\beta}\prod_{x\in\mathcal{D}}\sum_{k}[f_{k}(x;\theta)]^{\beta}.$
@@ 3.2 Supporting theory
The proof can be found in Appendix E. As our goal is to select $\beta$ that maximizes test LPD, a reasonable strategy is to optimize the lower bound presented in Lemma 3.1 with respect to $\beta$. As the second term in the lower bound is independent of $\beta$, this is therefore equivalent to maximizing the first term, $\mathbb{E}_{q(x,y)}\log p(y|x,\hat{\theta}_{\text{MAP}},\beta)$. Note that this objective function requires $\hat{\theta}_{\text{MAP}}$, which is often unavailable in closed-form. A straightforward solution is to replace it with an estimate from an iterative optimizer, before optimizing again with respect to $\beta$. This comes at the cost of two optimization runs (and the effort to tune their hyperparameters). Instead, we propose maximizing the empirical version of $\mathbb{E}_{q(x,y)}\log p(y|x,\theta,\beta)$ with respect to $\theta$ and $\beta$ simultaneously, leading us back to (7). We study the efficacy of this method empirically in the next section.
@@ 4 Experiments
We now illustrate the behavior of the PPDs, (3) and (8), across different values of $\beta$ on a suite of benchmark datasets for both regression and classification tasks. We refer to (3) as SM-PD and (8) as TM-PD, where SM and TM stand for standard model and tempered model, respectively. Notably, our data-driven procedure for selecting $\beta$ is agnostic to its downstream use in either SM-PD or TM-PD. We demonstrate that the proposed method performs well in both according to the test LPD metric. While our method aims to maximize test LPD and is supported by the theory discussed in Section 3.2, we also evaluate the point prediction of the PPD. Specifically, we use $\hat{y}=\mathbb{E}_{p_{\beta}(y|x,\mathcal{D})}[y]$ to compute mean squared error (MSE) $\frac{1}{|\mathcal{D_{\mathrm{test}}}|}\sum_{(x,y)\in\mathcal{D_{\mathrm{test}}}}(y-\hat{y})^{2}$ for regression tasks, and $\hat{y}=\arg\max_{y}p_{\beta}(y|x,\mathcal{D})$ to compute accuracy $\frac{1}{|\mathcal{D_{\mathrm{test}}}|}\sum_{(x,y)\in\mathcal{D_{\mathrm{test}}}}\mathds{1}[y=\hat{y}]$ for classification tasks. By definition, these point predictions $\hat{y}$ are identical across SM-PD and TM-PD, and their results are consolidated in our reporting. We compare our method against a grid search over $\beta\in\{0.1,0.3,1,3,\ldots,1000\}$, corresponding to increments of roughly 0.5 on the $\log_{10}$ scale. The optimal $\beta$ from the grid search is selected based on test LPD, MSE, or accuracy on a validation set [^2] In contrast, our method tracks the log-likelihood of $p(y|x,\theta,\beta)$ on a validation set, and the optimal $\beta$ may differ depending on the metric used.

### v2

@@ Abstract
The Cold Posterior Effect (CPE) is a phenomenon in Bayesian Deep Learning (BDL), where tempering the posterior to a cold temperature often improves the predictive performance of the posterior predictive distribution (PPD). Although the term ‘CPE’ suggests colder temperatures are inherently better, the BDL community increasingly recognizes that this is not always the case. Despite this, there remains no systematic method for finding the optimal temperature beyond grid search. In this work, we propose a data-driven approach to select the temperature that maximizes test log-predictive density, treating the temperature as a model parameter and estimating it directly from the data. We empirically demonstrate that our method performs comparably to grid search, at a fraction of the cost, across both regression and classification tasks. Finally, we highlight the differing perspectives on CPE between the BDL and Generalized Bayes communities: while the former primarily emphasizes the predictive performance of the PPD, the latter prioritizes the utility of the posterior under model misspecification; these distinct objectives lead to different temperature preferences.
@@ 1 Introduction
Recent decades have seen substantial advances in machine learning, with deployments across diverse applications. These range from image recognition and natural language processing to autonomous vehicles and medical diagnostics. Driven in no small part by the technological advances that enables large-scale training of high-fidelity models, these models obtain impressive generalization performance that are not readily explained with classical statistical wisdom.
@@
There is growing recognition in the BDL community that, despite the term ‘CPE’, colder temperatures do not always result in better predictive performance for the PPD (Adlam et al., 2020; Zhang et al., 2024). Unfortunately, the common approach of temperature tuning via grid search is computationally expensive, as it requires extra posterior sampling for multiple $T$ across the grid. To address these issues, we propose a data-driven method to select an appropriate temperature. To the best of our knowledge, no such dedicated tool exists without appealing to intermediate approximations (e.g., variational inference (Laves et al., 2021)). Our method only requires maximizing a likelihood function to find a suitable temperature. This can usually be done as part of the sampler warm-up phase, and importantly, does not require any extra posterior sampling.
@@ 2 Tempered Posteriors and the Cold Posterior Effect
$p_{\beta}(y|x,\mathcal{D})=\int p(y|x,\theta)p_{\beta}(\theta|\mathcal{D})\,\textrm{d}\theta,$ (3)
which is constructed from Bayesian model averaging, can achieve better performance (in terms of the test log predictive density defined below) in regression and classification by artificially tempering the posterior to $\beta>1$ (Wenzel et al., 2020). While our construction in (2) follows the convention of tempering both the likelihood and the prior (Wenzel et al., 2020; Fortuin et al., 2022), improvements in performance have also been observed when only tempering the likelihood (Aitchison, 2021; Bachmann et al., 2022; Kapoor et al., 2022).
@@
where the hat over $p$ indicates dependence on the training data $\mathcal{D}$, and the expectation is only taken over the ‘new’ observation $(x,y)$. Note that the test LPD is sometimes referred to as the ‘test log-likelihood’ in the literature, and the negative test LPD is often called ‘negative log-likelihood (NLL)’. However, we will avoid these terms as they can sound ambiguous. It can be shown that a predictive density $\hat{p}(y|x)$ with a higher test LPD is closer to the truth $q(y|x)$ in Kullback-Leibler (KL) divergence $\mathbb{E}_{q(x)}D_{\mathrm{KL}}(q(y|x)\|\hat{p}(y|x))$. Our primary interest in this work is to select $\beta$ such that the test LPD of (3) is high.
@@ 2.1 Outline
Following the reasoning that the aleatoric uncertainty in the data can be quantified by $\beta$ (Adlam et al., 2020; Kapoor et al., 2022), we advocate a likelihood-based approach to select $\beta$ directly from the data; details are presented in Section 3. We empirically show that our method can select a $\beta$ that is near-optimal in terms of test LPD. Experimental results are presented in Section 4.
We then conclude with a detailed discussion to dispel some misconceptions surrounding the CPE. In particular, we review recent work in BDL in Section 5, and explains why colder temperatures do not give better test LPD. We also address the converse hypothesis sometimes encountered in Generalized Bayes (GB) — that warmer temperatures are often better — in Section 6. This hypothesis may hold in certain contexts, e.g., ensuring posterior consistency (Grünwald, 2012), but does not usually improve test LPD.
We emphasize that these two communities prioritize different objectives: in BDL, the primary focus is on the PPD $p(y|x,\mathcal{D})$ (e.g., reliable uncertainty quantification of $y$), while in GB, the emphasis lies on the utility of the posterior $p(\theta|\mathcal{D})$ (e.g., fast concentration rate as $n\to\infty$, or credible intervals of $\theta$ with good coverage). Furthermore, the types of statistical models and datasets explored in each community are notably distinct. In BDL, large-scale neural networks are trained on extensive, nearly noiseless datasets, whereas in GB, the datasets are typically smaller, noisier, and the models are classical regular statistical models. Consequently, it is not surprising that each community arrives at different recommendations regarding temperature tuning.
@@ 3 Temperature Selection for Test LPD
It is commonly believed that the PPD (3) at $\beta=1$ is optimal for test LPD when the model is well-specified (Adlam et al., 2020; Aitchison, 2021). However, this is not always the case (see Appendix A for a counter-example). In fact, in the likelihood tempering case, Zhang et al. (2024, Theorem 4) shows that $\beta=1$ can only be optimal if and only if the training loss, $\sum_{(x,y)\in\mathcal{D}}\log p(y|x,\mathcal{D})$ remains unchanged with the inclusion of new data. We arrived at a similar conclusion for posterior tempering; see Appendix B for details. This motivates the development of an efficient method to select the optimal $\beta$.
3.1 On the Role of Temperature in PPD via Singular Learning Theory
If we could have some theoretical grasp on how temperature affects the test LPD of the PPD in (3), that is, $\operatorname{LPD}(p_{\beta}(y|x,\mathcal{D}))$, then it might suggest a methodology for temperature selection. A little known result to the BDL and GB communities is the following from singular learning theory (Watanabe, 2010a, Lemma 3), which applies to both regular and singular models $p(y|x,\theta)$  [^1] A model $p(y|x,\theta)$ is regular if the corresponding Fisher information matrix, $\mathbb{E}_{q(x,y)}[\nabla_{\theta}\log p(y|x,\theta)(\nabla_{\theta}\log p(y|x,\theta))^{\top}]$, is positive-definite. Otherwise, the model is singular.:
$-\mathbb{E}_{\mathcal{D}}\operatorname{LPD}(p_{\beta}(y|x,\mathcal{D}))$
$\quad=-\mathbb{E}_{\mathcal{D}}\mathbb{E}_{q(x,y)}\log p_{\beta}(y|x,\mathcal{D})$
$\quad=-\mathbb{E}_{q(x,y)}\log p(y|x,\theta_{\dagger})+\left[\frac{\lambda-\nu(\beta)}{\beta}+\nu(\beta)\right]\frac{1}{n}$
$\qquad+o\left(\frac{1}{n}\right)$ (4)
@@ 3.1 On the Role of Temperature in PPD via Singular Learning Theory
Some comments on (4) are in order. This result allows for both misspecification ($\theta_{\dagger}$ is not necessarily such that the KL divergence between the truth and $p(y|x,\theta_{\dagger})$ is zero) and singular models $p(y|x,\theta)$, including neural networks, as well as classical models satisfying standard regularity conditions. On the other hand, it is an asymptotic result in the sample size $n$, and hence the prior plays no role. We also remark that the relation in (4) pertains to the average (negative) test LPD of the PPD in (3), with the average taken over the training set $\mathcal{D}$, as indicated by the notation $\mathbb{E}_{\mathcal{D}}$. More precise interpretation of (4) can be divided into two settings:
@@
Let $d$ be the dimension of $\theta$. For well-specified regular models, $\lambda=\nu(\beta)=d/2$ for all $\beta$ (Watanabe, 2009), and the second term, $\left[\frac{\lambda-\nu(\beta)}{\beta}+\nu(\beta)\right]\frac{1}{n}$, reduces to $\frac{d}{2n}$. This implies that for large $n$ relative to $d$, temperature has little impact (though it may appear in higher-order terms in the expansion). This aligns with recent work by McLatchie et al. (2024), which arrives at a similar conclusion using different techniques. Under model misspecification, the second term can alternatively be expressed as $\frac{\beta}{n}\mathbb{E}_{\mathcal{D}}V(n)$ (Watanabe, 2010b), where $V(n)$ represents the functional variance and now depends on the temperature in a non-trivial manner.
@@ 3.2 Selecting Temperature Using the Tempered Model
In this section, we introduce a method to select $\beta$ that corresponds to high test LPD in both regression and classification tasks. Our approach is grounded in the insight that the tempered posterior can be reverse-engineered as the posterior from an alternative model-prior pair (Zeno et al., 2020; Zhang et al., 2024). Following these works, we define the tempered model,
@@
where the ‘rest of the terms’, $\tilde{p}(\theta|\beta)\propto p(\theta)^{\beta}\prod_{x\in\mathcal{D}}\int p(y^{\prime}|x,\theta)^{\beta}\textrm{d}y^{\prime}$, can be viewed as a prior on $\theta$ with the ‘normalizing constant’ being a function of $x$ in $\mathcal{D}$. The term $\tilde{p}(\theta|\beta)$ can be seen as an input-dependent prior (Zeno et al., 2020) but we suppress the dependence on $x$ in the notation. The (inverse) temperature $\beta$ has an intuitive interpretation of controlling the ‘spikiness’ of the tempered model. We give two examples with Gaussian (regression) and categorical (classification) models:
@@
$p(y|x,\theta,\beta)=\mathcal{N}\left(y\left|\,\mu(x;\theta),\frac{\sigma^{2}}{\beta}\right.\right),$
$\tilde{p}(\theta|\beta)\propto p(\theta)^{\beta}\left(\dfrac{2\pi\sigma^{2}}{\beta(2\pi\sigma^{2})^{\beta}}\right)^{n/2}.$
@@
$p(y|x,\theta,\beta)=\dfrac{\exp(\beta f_{y}(x;\theta))}{\sum^{K}_{k=1}\exp(\beta f_{k}(x;\theta))},$
$\tilde{p}(\theta|\beta)\propto p(\theta)^{\beta}\prod_{x\in\mathcal{D}}\sum_{k}[f_{k}(x;\theta)]^{\beta}.$
@@
A similar temperature optimization approach was proposed in Guo et al. (2017) for computing a well-calibrated ‘plug-in’ predictive density $p(y|x,\theta^{*}_{\text{SGD}},\beta)$, where $\theta^{*}_{\text{SGD}}$ is typically an SGD solution of neural networks from a standard training workflow. In their method, the optimal $\beta$ is computed post hoc by maximizing $\frac{1}{n}\sum_{(x,y)\in\mathcal{D_{\mathrm{valid}}}}p(y|x,\theta^{*}_{\text{SGD}},\beta)$ on a validation set, with $\theta$ fixed and only $\beta$ being optimized. While the difference appears subtle, we find that jointly optimizing $\theta$ and $\beta$ is the key to a good $\beta$ for constructing PPDs. Detailed experimental results are provided in Appendix D.
@@ 3.3 Supporting Theory
$\operatorname{LPD}(\mathbb{E}_{p_{\beta}(\theta|\mathcal{D})}[p(y|x,\theta,\beta)])$
$\quad>\mathbb{E}_{q(x,y)}\log p(y|x,\hat{\theta}_{\textrm{MAP}},\beta)-\frac{1}{2}\mathbb{E}_{q(x,y)}\log{(1+x^{\top}\bm{\Sigma}x)},$
@@
The proof can be found in Appendix E. As our goal is to select $\beta$ that maximizes test LPD, a reasonable strategy is to optimize the lower bound presented in Lemma 3.1 with respect to $\beta$. As the second term in the lower bound is independent of $\beta$, this is therefore equivalent to maximizing the first term, $\mathbb{E}_{q(x,y)}\log p(y|x,\hat{\theta}_{\text{MAP}},\beta)$. However, this objective function requires $\hat{\theta}_{\text{MAP}}$, which is often unavailable in closed-form. A straightforward solution is to replace it with an estimate from an iterative optimizer, before optimizing again with respect to $\beta$. This comes at the cost of two optimization runs (and the effort to tune their hyperparameters). Instead, we propose maximizing the empirical version of $\mathbb{E}_{q(x,y)}\log p(y|x,\theta,\beta)$ with respect to $\theta$ and $\beta$ simultaneously, leading us back to (7). We study the efficacy of this method empirically in the next section.
@@ 4 Experiments
We now illustrate the behavior of the PPDs, (3) and (8), across different values of $\beta$ on a suite of benchmark datasets for both regression and classification tasks. We refer to (3) as SM-PD and (8) as TM-PD, where SM and TM stand for standard model and tempered model, respectively. Notably, our data-driven procedure for selecting $\beta$ is agnostic to its downstream use in either SM-PD or TM-PD. We demonstrate that the proposed method performs well in both in terms of test LPD. While our method aims to maximize test LPD and is supported by the theory discussed in Section 3.3, we also evaluate the point predictions of the PPD. Specifically, we use $\hat{y}=\mathbb{E}_{p_{\beta}(y|x,\mathcal{D})}[y]$ to compute mean squared error (MSE) $\frac{1}{|\mathcal{D_{\mathrm{test}}}|}\sum_{(x,y)\in\mathcal{D_{\mathrm{test}}}}(y-\hat{y})^{2}$ for regression, and $\hat{y}=\arg\max_{y}p_{\beta}(y|x,\mathcal{D})$ to compute accuracy $\frac{1}{|\mathcal{D_{\mathrm{test}}}|}\sum_{(x,y)\in\mathcal{D_{\mathrm{test}}}}\mathds{1}[y=\hat{y}]$ for classification. By definition, these point predictions $\hat{y}$ are identical across SM-PD and TM-PD, and their results are consolidated in our reporting. We compare our method against a grid search over nine $\beta\in\{0.1,0.3,1,3,\ldots,1000\}$, corresponding to increments of roughly 0.5 on the $\log_{10}$ scale. The optimal $\beta$ from the grid search is selected based on test LPD, MSE, or accuracy on a validation set [^2] In contrast, our method tracks the log-likelihood of $p(y|x,\theta,\beta)$ on a validation set, and the optimal $\beta$ may differ depending on the metric used.